Grade upto college level Physical Chemistry

An insulated container contains 1 mol of a liquid, molar volume 100 ml, at 1 bar. When liquid is steeply pressed to 100 bar, volume decreases to 99 ml. Find. ∆H and ∆U for the process.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
For adiabatic process, W = P (V2 – V1)
Here P1 = 1 bar, P2= 100 bar, V1 = 100 mL, V2 = 99 mL; For adiabatic process, q = 0 / ∆U = w
∆U = q + W
= q – P(V2 – V1) since W = - P(V2= - V1)
= 0 – 100 (99 - 100) = 100 bar mL
∆H = ∆U + ∆(PV) = ∆U + (P2V2 – P1V1)
100 + [(100 * 99) – (1 * 100)]
= 100 + (9900 - 100) = 9900 bar mL

Thanks
Deepak patra
askIITians Faculty